### v1

# Hongzhi Zang1, Mingjie Wei6,2, Si Xu3, Yongji Wu5, Zhen Guo3, Yuanqing Wang4,3, Hao Lin3, &Liangzhi Shi1, Yuqing Xie1, Zhexuan Xu1, Zhihao Liu7,2, Kang Chen4,2, Wenhao Tang1, &Quanlu Zhang3, Weinan Zhang6, Chao Yu1,2,†,∗, Yu Wang1,†

Abstract
Recent progress in vision and language foundation models has significantly advanced multimodal understanding, reasoning, and generation, inspiring a surge of interest in extending such capabilities to embodied settings through vision-language-action (VLA) models. Yet, most VLA models are still trained with supervised fine-tuning (SFT), which struggles to generalize under distribution shifts due to error accumulation. Reinforcement learning (RL) offers a promising alternative by directly optimizing task performance through interaction, but existing attempts remain fragmented and lack a unified platform for fair and systematic comparison across model architectures and algorithmic designs. To address this gap, we introduce RLinf-VLA, a unified and efficient framework for scalable RL training of VLA models. The system adopts a highly flexible resource allocation design that addresses the challenge of integrating rendering, training, and inference in RL+VLA training. In particular, for GPU-parallelized simulators, RLinf-VLA implements a novel hybrid fine-grained pipeline allocation mode, achieving a 1.61x-1.88x speedup in training. Through a unified interface, RLinf-VLA seamlessly supports diverse VLA architectures (e.g., OpenVLA, OpenVLA-OFT), multiple RL algorithms (e.g., PPO, GRPO), and various simulators (e.g., ManiSkill, LIBERO). In simulation, a unified model achieves 98.11% across 130 LIBERO tasks and 97.66% across 25 ManiSkill tasks. Beyond empirical performance, our study distills a set of best practices for applying RL to VLA training and sheds light on emerging patterns in this integration. Furthermore, we present preliminary deployment on a real-world Franka robot, where RL-trained policies exhibit stronger generalization than those trained with SFT. We envision RLinf-VLA as a foundation to accelerate and standardize research on embodied intelligence.
Figure 1: Overview of RLinf-VLA. Through a unified interface, RLinf-VLA seamlessly supports diverse VLA architectures, multiple RL algorithms, and various simulators. To flexibly accommodate the integration of rendering, training, and inference in RL+VLA training, RLinf-VLA provides three GPU allocation modes: colocated, disaggregated, and a novel hybrid mode. A single unified model achieves 98.11% success on 130 LIBERO tasks and 97.66% on 25 ManiSkill tasks in simulation, while an RL-trained model demonstrates superior zero-shot generalization on a real-world Franka robot compared to the SFT model.
1 Introduction
Vision-Language-Action (VLA) models represent a new generation of foundation models that aim to unify perception, language understanding, and embodied control (Ma et al., 2024; Firoozi et al., 2025). By leveraging vision-language models pretrained on internet-scale data and further training them on large, heterogeneous robot demonstration datasets (O’Neill et al., 2024; Khazatsky et al., 2024), VLAs can map raw sensor observations and natural language instructions directly into robot actions. This paradigm has already shown strong generalization across a wide range of tasks (Team et al., 2024; Kim et al., 2024; Liu et al., 2024; Wen et al., 2025; Shah et al., 2023; Black et al., 2024; Intelligence et al., 2025). However, the dominant training paradigm for VLAs remains supervised fine-tuning (SFT) via behavioral cloning on demonstration data (Kim et al., 2025). While effective in matching expert trajectories, SFT is highly vulnerable to distribution shift: small deviations from expert behavior can compound over time, pushing the policy into unfamiliar states and degrading robustness (Ross & Bagnell, 2010; De Haan et al., 2019; Belkhale et al., 2023; Foster et al., 2024). This mismatch between training and deployment distributions fundamentally limits scalability and generalization.
Reinforcement learning (RL) offers a complementary solution by directly optimizing cumulative task rewards through trial and error (Sutton & Barto, 1998). Unlike SFT, RL enables exploration beyond narrow expert data and equips policies with corrective strategies. Recent evidence suggests that RL fine-tuning can yield stronger out-of-distribution performance than SFT, particularly in both semantic alignment and execution robustness (Liu et al., 2025). Yet, progress on applying RL to VLAs remains small-scale (Liu et al., 2025) or fragmented (Lu et al., 2025; Li et al., 2025), primarily due to the lack of a unified and efficient framework to support large-scale experimentation. Unlike RL post-training for LLMs or VLMs, RL for VLAs requires repeated interactions with environments. These environments are typically realized through simulators, which also consume significant GPU memory and thus compete with model training and inference for memory resources—an issue that existing large-scale RL frameworks handle inefficiently. In addition, different works adopt diverse models, algorithms, and evaluation setups, making it difficult to compare methods fairly or distill general principles.
To address these challenges, we present RLinf-VLA, a unified and efficient framework for scalable RL training of VLA models.
  • Unified: RLinf-VLA provides a cohesive interface that brings together multiple dimensions of RL for VLA training. It integrates different simulators, including ManiSkill (Tao et al., 2025) and LIBERO (Liu et al., 2023a), with features such as partial environment reset. It supports multiple algorithms, implementing both Proximal Policy Optimization (PPO) (Schulman et al., 2017; Ouyang et al., 2022) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Guo et al., 2025). The framework is compatible with diverse VLA architectures, including OpenVLA (Kim et al., 2024) and OpenVLA-OFT (Kim et al., 2025) with action chunking. To flexibly accommodate the integration of rendering, training, and inference in RL+VLA training, RLinf-VLA provides three GPU allocation modes: colocated, disaggregated, and a novel hybrid mode. Switching between these modes can be achieved through straightforward configuration adjustments.
  • Efficient: Beyond unification, RLinf-VLA is designed for efficiency at both the system and algorithmic levels. It achieves a 2.27× improvement in throughput compared to the baseline framework. In particular, within GPU-parallelized simulators, the hybrid fine-grained pipelining mechanism further accelerates RL training, yielding a 1.61×–1.88× speedup. On the algorithmic side, RLinf-VLA incorporates several design choices, including lightweight critics, loss normalization, action masking, and rollout filtering, which collectively substantially enhance training efficiency.
Extensive experiments validate the effectiveness of RLinf-VLA. In simulation, a single unified model attains 98.11% success across 130 LIBERO tasks for the first time and 97.66% success across 25 ManiSkill tasks. More importantly, our study distills a set of best practices for RL-based VLA training:
  • PPO: action-level value estimation is superior to chunk-level estimation when using action chunks, and partial resets substantially improve sample efficiency.
  • GRPO: trajectory-length normalization and action masking are essential for stable training.
  • General Practices: larger rollout batch size is useful; LoRA may not directly affect performance but typically requires hyperparameter re-tuning.
Finally, we demonstrate preliminary deployment on a real-world Franka robot, where RL-trained policies exhibit stronger zero-shot generalization than those trained with SFT. Specifically, in a pick-and-place task with six previously unseen objects, the RL policy successfully completed 8 out of 30 trials, whereas the SFT policy failed in all 30 trials under the same zero-shot conditions. RLinf-VLA has been open-sourced and is maintained with frequent updates. We envision RLinf-VLA as a foundation to accelerate, standardize, and scale research in embodied intelligence.
Our contribution can be summarized as follows:
  • Unified design: RLinf-VLA supports multiple simulators (ManiSkill, LIBERO), diverse VLA architectures (OpenVLA, OpenVLA-OFT), and RL algorithms (PPO, GRPO), with three execution modes, including a novel hybrid GPU allocation mode, enabling scalable and configurable training.
  • Efficiency in system and algorithm design: Hybrid fine-grained pipelining for GPU-parallelized simulators, colocated allocation for CPU-parallelized simulators, and algorithmic enhancements such as lightweight critics, loss normalization, action masking, and rollout filtering substantially improve training efficiency.
  • Strong empirical performance and generalization: A single unified model achieves 98.11% success on 130 LIBERO tasks and 97.66% on 25 ManiSkill tasks in simulation, while an RL-trained model demonstrates superior zero-shot generalization on a real-world Franka robot compared to the SFT model.
  • Open-source and maintainable platform: RLinf-VLA is publicly released with frequent updates, providing a foundation to accelerate, standardize, and scale research in embodied intelligence.
2 Preliminary for Reinforcement Learning
In this paper, we finetune VLA models with reinforcement learning to solve vision-based manipulation tasks, which we formulate as a partially observable Markov decision process (POMDP). A POMDP extends the standard MDP and is defined by the tuple $(\mathcal{S},\mathcal{A},P,r,\gamma,\Omega,O)$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space (discrete or continuous, potentially chunked), $P(s^{\prime}\mid s,a)$ the state transition function, $r(s,a)$ the reward function, $\gamma\in[0,1]$ the discount factor, $\Omega$ the observation space, and $O(o\mid s)$ the observation function mapping a state $s$ to an observation $o$.
At each timestep $t$, the agent samples an action $a_{t}\sim\pi_{\theta}(\cdot\mid o_{t})$ based on the current observation $o_{t}\sim O(\cdot\mid s_{t})$. The environment then executes $a_{t}$, transitions to the next state $s_{t+1}$ via $P$, and produces the next observation $o_{t+1}$. The goal of RL is to optimize the parameterized policy $\pi_{\theta}$ so as to maximize the expected cumulative discounted reward over trajectories $\tau=(s_{0},a_{0},\dots,s_{T})$:
$J(\theta)=\mathbb{E}_{\tau\sim\pi_{\theta}}\Bigg[\sum_{t=0}^{T}\gamma^{t}r(s_{t},a_{t})\Bigg].$ (1)
We now recall several standard definitions in RL:
Definition 1 (Return).
The return $R_{t}$ is the $\gamma$-discounted cumulative reward starting from timestep $t$:
$R_{t}=\sum_{k=0}^{\infty}\gamma^{k}r(s_{t+k},a_{t+k}).$ (2)
Definition 2 (Value Function).
The value function $V_{\pi}(s)$ is the expected return obtained from state $s$ at timestep $t$ when following policy $\pi$:
$V_{\pi}(s)=\mathbb{E}_{\pi}[R_{t}\mid s_{t}=s].$ (3)
Definition 3 (Action-Value Function).
The action-value function $Q_{\pi}(s,a)$ is the expected return obtained by executing action $a$ in state $s$ at timestep $t$ and thereafter following policy $\pi$:
$Q_{\pi}(s,a)=\mathbb{E}_{\pi}[R_{t}\mid s_{t}=s,a_{t}=a].$ (4)
Definition 4 (Advantage Function).
The advantage function $A_{\pi}(s,a)$ quantifies how much better taking action $a$ in state $s$ is compared to the expected value of the state:
$A_{\pi}(s,a)=Q_{\pi}(s,a)-V_{\pi}(s).$ (5)
3 Framework Design
3.1 GPU Allocation Strategies
Figure 2: VLA+RL training loop
In robotic reinforcement learning, GPU resources are primarily consumed by the training process (Training) and the rollout process (Generation and Simulator). The training loop is shown in Figure 2. Simulators, however, vary by type: CPU-parallelized and GPU-parallelized simulators exhibit fundamentally different patterns of resource utilization. In CPU-parallelized simulators, the simulation logic is distributed across multiple CPU workers, while the GPU is mainly utilized for rendering and policy inference. As a result, CPU cores often become the primary bottleneck limiting overall performance. By contrast, GPU-parallelized simulators integrate rendering, physics simulation, and inference directly on the GPU, exploiting massive parallelism to achieve significantly higher throughput. However, this tight coupling also causes different components to compete for GPU memory and compute resources, making the overall performance highly sensitive to resource scheduling. Consequently, CPU-parallelized and GPU-parallelized simulators require distinct GPU allocation strategies, and efficient, flexible allocation across simulation, training, and inference is essential to maximize throughput.
(a) Colocated GPU Allocation
(b) Disaggregated GPU Allocation
(c) Hybrid GPU Allocation with Fine-Grained Pipelining
Figure 3: Different GPU allocation strategies.
To address this issue, our framework supports three allocation modes: colocated, disaggregated, and hybrid. Users can easily switch between these modes by simply modifying a YAML configuration file, greatly reducing the overhead of system customization. In particular, for GPU-parallelized simulators, we introduce a novel hybrid allocation with fine-grained pipelining strategy that further improves utilization efficiency. Figure 3 illustrates the core ideas of these allocation modes, while their algorithmic features are summarized here. For the full system design of RLinf, we refer the reader to Yu et al. (2025a).
3.1.1 Colocated Allocation
In this mode, all components (Training, Generation, and Simulator) share all GPUs simultaneously. This maximizes data parallelism and ensures that no GPU is left completely idle. Figure 3(a) provides an overview of the colocated allocation mode.
Offload Support
To make colocated allocation feasible under memory constraints, we introduce offloading support. Idle components can be temporarily moved to CPU memory, freeing GPU resources for the active ones. For Training and Generation, this offload–onload process is relatively straightforward. For the Simulator, however, we must manually store all essential environment states, so that the simulator can resume correctly once reloaded to the GPU.
Despite this support, colocated allocation is not always efficient. Two problematic cases often arise: 1) When multiple components reside on the same GPU, they may occupy memory concurrently even though only one is actively computing, leading to wasted capacity. 2) Although offloading reduces memory contention, its overhead can become substantial in practice. In particular, the frequent interaction between Generation and Simulator during rollout makes repeated offload–onload operations prohibitively expensive.
3.1.2 Disaggregated Allocation
Figure 3(b) depicts the central idea of the disaggregated allocation mode. In this mode, each component is assigned to a distinct (possibly multi-GPU) partition, with no overlap across components. This ensures that every component can fully exploit its allocated resources. However, disaggregated allocation may still lead to GPU underutilization due to inter-component dependencies. For instance, the Simulator must wait for actions generated by the Generation, leaving some GPUs idle and creating “GPU bubbles.”
3.1.3 Hybrid Allocation with Fine-grained Pipelining
To overcome the drawbacks of the above modes, we propose a hybrid strategy: hybrid allocation combined with fine-grained pipelining. In hybrid allocation, components can flexibly select GPUs. A typical configuration is to assign Generation and Simulator to different GPU partitions, while allowing Training to utilize all GPUs.
On top of this, we introduce fine-grained pipelining to mitigate bubbles caused by inter-component dependencies. Specifically, a simulator instance on one GPU is partitioned into multiple sub-simulators, denoted as $S^{(1)},S^{(2)},\ldots,S^{(k)}$. The pipeline proceeds as follows [^1] In this section, superscripts denote simulator indices, while subscripts indicate timesteps.:
  1. 1.
    At step $t=0$, $S^{(1)}$ generates the initial observation $o_{0}^{(1)}$, which is sent to the Generation component to produce action $a_{0}^{(1)}$.
  2. 2.
    Meanwhile, $S^{(2)}$ generates $o_{0}^{(2)}$ in parallel.
  3. 3.
    Once $a^{(1)}_{0}$ is ready, it is fed back into $S^{(1)}$ to produce the next observation $o_{1}^{(1)}$, while $o_{0}^{(2)}$ is simultaneously processed by the actor to generate $a^{(2)}_{0}$.
This scheduling allows Simulator and Generation to run concurrently, reducing idle time while preserving correctness. In this way, our framework avoids the frequent offloading overhead of colocated allocation while also eliminating the GPU bubbles that occur in disaggregated allocation. Figure 3(c) provides a schematic illustration of the hybrid allocation mode with fine-grained pipelining ($k=2$).
3.1.4 Configuring GPU Allocation Modes
Our framework exposes a simple configuration interface that allows users to flexibly choose GPU allocation strategies without explicitly specifying the mode. Instead, users only need to assign GPU IDs for each component via the following fields:
  • cluster.component_placement.env for the Simulator,
  • cluster.component_placement.rollout for Generation,
  • cluster.component_placement.actor for Train.
In addition, offloading can be enabled or disabled independently for each component using:
  • env.enable_offload,
  • rollout.enable_offload,
  • actor.enable_offload.
Finally, fine-grained pipelining is controlled by the parameter rollout.pipeline_stage_num (corresponding to $k$ in Section 3.1.3). A value greater than 1 splits each simulator instance on a GPU into multiple pipeline stages, while a value of 1 disables pipelining.
3.2 Model Compatibility
Our framework is designed to be flexible and extensible, supporting both parameter-efficient training methods and different VLA model variants. In particular, we provide built-in support for Low-Rank Adaptation (LoRA) (Hu et al., 2021), as well as for the OpenVLA (Kim et al., 2024) and OpenVLA-OFT (Kim et al., 2025) models.
3.2.1 LoRA Support
LoRA (Hu et al., 2021) is a parameter-efficient fine-tuning method that adapts large pre-trained models to new tasks without updating all of their parameters. Instead of modifying the full weight matrices, LoRA introduces additional low-rank decomposition layers into linear transformations and trains only these small sets of parameters while keeping the original weights frozen. This greatly reduces the number of trainable parameters, lowering memory consumption and speeding up training while maintaining performance comparable to full fine-tuning. In our framework, users can flexibly choose whether to enable LoRA. If “is_lora” is enabled, all linear modules in VLA are designated as target modules by default, ensuring broad coverage of trainable parameters while preserving efficiency.
3.2.2 Model type
In RLinf-VLA, the argument “model_name” specifies the VLA model to be used, with current support for openvla” and “openvla_oft”. We briefly introduce these two models below.
OpenVLA
OpenVLA (Kim et al., 2024) is a VLA model containing about 7B parameters. Despite being smaller than some closed-source counterparts such as RT-2-X (55B) (O’Neill et al., 2024), OpenVLA demonstrates superior performance on a range of robotic manipulation benchmarks, achieving higher success rates while remaining computationally efficient. Our framework supports training and evaluation of OpenVLA, with extensive experiments conducted in the ManiSkill environment.
OpenVLA-OFT
OpenVLA-OFT (Optimized Fine-Tuning) (Kim et al., 2025) extends OpenVLA by focusing on fine-tuning efficiency, inference speed, and practical deployability in robotics. Instead of relying on discrete action tokens and complex objectives, OpenVLA-OFT represents actions in a continuous space and optimizes them using a simple L1 regression loss. This design reduces training overhead and accelerates convergence. Moreover, OpenVLA-OFT incorporates parallel decoding together with action chunking, which significantly improves inference throughput and enables real-time control at higher frequencies. These advances yield strong performance gains in both the LIBERO benchmark and real-world tasks, while also achieving over an order-of-magnitude speedup in action generation. Our framework supports training OpenVLA-OFT with a discretized action mode and provides experimental results on ManiSkill and LIBERO environments.
3.2.3 Unified Interface
3.3 Multiple Simulators Support
Simulators play a crucial role in reinforcement learning for robotics, providing controllable and efficient environments in which agents can interact, learn, and improve without the high costs and risks of physical hardware. Compared to real-world experiments, simulators offer several key advantages:
  • Scalability of data collection: By leveraging parallel computation, simulators can generate a large volume of interaction data (often at least thousands of frames per second), far exceeding what is feasible with physical robots.
  • Task diversity: Simulators provide access to a wide range of robotic tasks, including table-top manipulation, drawing/cleaning, and dexterous manipulation. This diversity is crucial for improving and evaluating the generalization capabilities of VLA models.
  • Safety and controllability: Agents can explore potentially dangerous or failure-prone behaviors without risking damage to hardware.
  • Fast environment reset and reduced human labor: Simulators enable instant and reproducible environment resets, greatly accelerating the training cycle and eliminating the need for manual intervention during repeated experiments.
Given these advantages, our framework focuses on implementing robotics RL in simulation. In this work, we select two representative platforms, ManiSkill and LIBERO, as our primary simulators. ManiSkill excels in physics-based manipulation tasks with diverse object assets, while LIBERO emphasizes instruction-conditioned tasks that test generalization and high-level reasoning for VLA models. Together, these simulators provide complementary capabilities, forming a solid foundation for training and evaluating our framework.
3.3.1 Unified Interface
Our framework supports vectorized environments with a unified interface across different simulators. For example, we leverage ManiSkill’s GPU-parallelized vectorized environments and LIBERO’s CPU-parallelized vector environment wrapper. On top of these backends, we define a consistent interface to seamlessly integrate with our training framework.
The unified interface consists of two categories: core functions and utility functions.
Core Functions
We implement standard Gym-style APIs, including “reset” and “step”. The “step” function additionally supports an “auto_reset” option: when enabled, any sub-environment that terminates or is truncated will be automatically reset, thereby improving sample efficiency by avoiding idle environments. Following the ManiSkill convention, we also support the “ignore_terminations” option. When enabled, termination signals are ignored and only truncation signals are respected, meaning that an episode ends only when the maximum episode length is reached. This feature allows us to flexibly support different implementation variants discussed in Section 4.2.1 and Section 4.2.2.
Beyond these, we extend the interface with a “chunk_step” function to handle action chunks. Instead of naively looping over the chunk with repeated “step” calls, this function manages episode termination more carefully. Two modes are supported: (1) reset immediately when a sub-environment finishes within the chunk, or (2) defer reset until the entire chunk has been executed. This flexibility ensures the correct handling of episode boundaries when chunked actions are used.
Utility Functions
Utility functions offer convenient support for training and evaluation. For instance, visualization utilities enable effortless generation of videos during rollouts or evaluation. In addition, we provide specialized utilities required by specific algorithms. For example, GRPO necessitates that all environments within a group share the same initial state, which can be ensured by setting “use_fixed_reset_state_ids=True”.
3.3.2 Tasks
We implement a variety of tasks based on the two simulators in our framework.
ManiSkill
Following RL4VLA (Liu et al., 2025), we construct the environment “PutOnPlateInScene25Main-v3”, which consists of 25 pick-and-place tasks with variations in objects and receptacles. We also adopt the out-of-distribution evaluation settings proposed in RL4VLA. By replicating their results, we further confirm the generalizability of VLA-RL models under these task variations.
LIBERO
We leverage the public benchmark task groups provided by LIBERO (Liu et al., 2023b), including LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-10, and LIBERO-90. Specifically:
  • LIBERO-Spatial: evaluates the model’s ability to reason about relative object positions.
  • LIBERO-Object: tests generalization across different object types.
  • LIBERO-Goal: examines the model’s capability to adapt to diverse task goals.
  • LIBERO-10: a long-horizon benchmark consisting of 10 tasks.
  • LIBERO-90: a long-horizon benchmark covering 90 tasks.
Most existing VLA+RL studies train and evaluate models within a single task group. In contrast, we train our model on the combined benchmark, LIBERO-130, which comprises 130 tasks specifically constructed to evaluate RLinf-VLA’s capability to learn under large-scale task settings, and we assess its performance across these task groups.
3.4 Multi-Algorithm Support
Our framework provides support for multiple reinforcement learning algorithms, with an initial focus on Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Guo et al., 2025). Both are representative on-policy algorithms that have been widely adopted in robotics research. PPO has established itself as a robust and general-purpose baseline for continuous control, while GRPO offers a simplified alternative by removing the need for value function estimation. Together, they form a solid foundation for training and evaluating VLA models in simulators.
3.4.1 PPO
PPO is one of the most widely adopted reinforcement learning algorithms in robotics. PPO enhances training stability by constraining policy updates within a trust region, thereby preventing overly large changes that could destabilize learning. This is achieved through a clipped surrogate objective, which balances exploration and exploitation while maintaining sample efficiency. Due to its robustness and simplicity, PPO has become a standard baseline in robotics RL and serves as the foundation of our framework.
In PPO, the advantage function is commonly estimated using Generalized Advantage Estimation (GAE) (Schulman et al., 2015):
$\hat{A}_{t}=\sum_{k=0}^{T-t-1}(\gamma\lambda)^{k}\big(r_{t+k}+\gamma V(s_{t+k+1})-V(s_{t+k})\big),$ (6)
where $r_{t}$ denotes the reward at timestep $t$, $V(s)$ is the value function, $\gamma$ is the discount factor, $\lambda$ controls the bias–variance trade-off, and $T$ is the episode horizon.
The PPO optimization objective is defined as:
$J^{\text{PPO}}(\theta)=\mathbb{E}_{t}\Big[\min\big(\rho_{t}(\theta)\hat{A}_{t},\;\mathrm{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\big)\Big],$ (7)
where
$\rho_{t}(\theta)=\frac{\pi_{\theta}(a_{t}\mid o_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid o_{t})},$ (8)
$\pi_{\theta_{\text{old}}}$ denotes the rollout policy, and $\epsilon$ is the clipping parameter. [^2] The definition of action $a_{t}$ differs across LLM and robotics RL settings. Here we treat it as a generalized action, with the distinctions discussed in Section 4.1.2. The same applies to GRPO.
3.4.2 GRPO
GRPO is a recent variant of policy optimization designed to simplify reinforcement learning pipelines. Unlike PPO, which requires training both a policy model and a value model, GRPO eliminates the need for an explicit value function by leveraging group-based relative comparisons of trajectories. This design reduces the overall model complexity and avoids potential inaccuracies from value estimation. As a result, GRPO provides a lightweight yet effective alternative to PPO, making it especially attractive for large-scale VLA training where efficiency and simplicity are important. We begin with a brief overview of GRPO as applied in LLM reasoning models, and present its specific design for robotics in Section 4.1.2. In GRPO, the important ratio can be defined as:
$\rho_{t}^{(i)}(\theta)=\frac{\pi_{\theta}(a_{t}^{(i)}\mid o_{t}^{(i)})}{\pi_{\theta_{\text{old}}}(a_{t}^{(i)}\mid o_{t}^{(i)})},$ (9)
where $i$ denotes $i$-th trajectory. But the advantage can be defined as:
$\hat{A}^{(i)}=\hat{A}_{t}^{(i)}=\frac{\mathcal{R}^{(i)}-\mathrm{mean}(\{\mathcal{R}^{(j)}\}_{j=1}^{G})}{\mathrm{std}(\{\mathcal{R}^{(j)}\}_{j=1}^{G})},\quad\forall t\in\{0,1,\ldots,|\tau^{(i)}|\},$ (10)
where $\mathcal{R}^{(i)}$ denotes the total reward of trajectory $\tau^{(i)}$. Note that $\mathcal{R}$ differs from the discounted return $R$, and $G$ represents the group size. Given an initial observation $o_{0}$ from sample buffer $D$, the behavior model $\pi_{\theta_{\text{old}}}$ generates $G$ trajectories $\{\tau_{i}\}_{i=1}^{G}$. The GRPO optimization objective is:
$J^{\text{GRPO}}(\theta)=\mathbb{E}_{o_{0}\sim D,\ \{\tau^{(i)}\}\sim\pi_{\theta_{\text{old}}}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|\tau^{(i)}|}\sum_{t=1}^{|\tau^{(i)}|}\min\Bigl(\rho_{t}^{(i)}(\theta)\,\hat{A}^{(i)},\;\mathrm{clip}\bigl(\rho_{t}^{(i)}(\theta),\,1-\epsilon,\,1+\epsilon\bigr)\,\hat{A}^{(i)}\Bigr)\Bigg],$ (11)
where $|\tau^{(i)}|$ is the length of trajectory $\tau^{(i)}$, $\epsilon$ is the clip parameter.
4 Algorithm Design
4.1 Advantage and Log-Probability
4.1.1 Advantage with Action Chunks
Figure 4: Illustration of different logprob levels.
Action Chunk
An action chunk refers to having the policy predict a short sequence of future actions rather than a single-step action at each timestep. This approach has been shown to better handle non-Markovianity in offline data, such as the non-Markovian behaviors often present in human demonstrations (Zhao et al., 2023). The technique is widely adopted in imitation learning. To integrate action chunks into reinforcement learning, we must define how advantages are associated with actions inside a chunk. Consider the $t$-th chunk $c_{t}=(a_{t,1},a_{t,2},\ldots,a_{t,C})$, where $C$ denotes the chunk length. Each atomic action $a_{t,j}$ produces an immediate reward $r_{t,j}$. Two formulations are commonly used:
  1. 1.
    Chunk-level: The entire chunk $c_{t}$ is treated as a macro-action, with the reward defined as the sum of per-step rewards, i.e., $r_{t}=\sum_{j=1}^{C}r_{t,j}$.
  2. 2.
    Action-level: Each $a_{t,j}$ is treated as an individual action, with its own reward and advantage.
These two formulations lead to different behaviors under GAE-based estimation.
Our framework supports both options. Users can control the advantage formulation via the parameter reward_type, which can be set to either chunk_level or action_level.
4.1.2 Log-Probability Granularity
In algorithms such as PPO and GRPO, computing the log-probabilities of actions under the current policy is essential for estimating the importance sampling ratios $\rho(\theta)$ in Equations (8) and (9). However, the precise definition of an “action” depends on the task domain.
In robotics RL, probabilities are typically computed either at the action level or the chunk level, depending on whether each environment step is considered atomic or grouped into a chunk. In contrast, in LLM-based reasoning tasks, policies generate sequences of tokens, where the natural atomic unit is a token. In this case, action probabilities are computed at the token level, by multiplying (or summing in log-space) the probabilities of individual tokens.
Thus, our framework supports three granularity choices for log-probability computation: action_level, chunk_level, and token_level, as shown in Figure 4. The selection is controlled by the parameter logprob_type, and each option leads to different optimization dynamics. Formally, we define a trajectory as $\tau=(c_{0},c_{1},\ldots,c_{T})$, where each chunk $c_{t}$ contains $C$ atomic actions, $c_{t}=(a_{t,1},a_{t,2},\ldots,a_{t,C})$, and each atomic action $a_{t,i}$ consists of $M$ tokens, $a_{t,i}=(d_{t,i,1},d_{t,i,2},\ldots,d_{t,i,M})$. The policy probabilities decompose as:
  • Chunk-level: $\pi(c_{t}|o_{t})=\prod_{i=1}^{C}\pi(a_{t,i}|o_{t},a_{t,:i-1})$,
  • Action-level: $\pi(a_{t,i}|o_{t},a_{t,:i-1})=\prod_{j=1}^{M}\pi(d_{t,i,j}|o_{t},d_{t,i,:j-1})$,
  • Token-level: $\pi(d_{t,i,j}|o_{t},d_{t,i,:j-1})$.
4.1.3 Supported Granularity Combinations for Advantage and Log-Probability
Table 1 summarizes the supported combinations of advantage and log-probability granularities. In general, an advantage type is compatible with all log-probability types whose granularity is no finer than its own. When the log-probability is defined at a finer granularity, we apply a broadcasting mechanism to the advantage values. For instance, if the advantage is defined at the chunk level while the log-probability is computed at the token level, the same chunk-level advantage is assigned to each token within the chunk.
Table 1: Supported combinations of advantage and log-probability granularities.
Advantage Log-Probability Chunk-level Action-level Token-level
Chunk-level
Action-level
4.2 Design Choices
4.2.1 PPO
Partial Reset Support
In reinforcement learning rollouts, a sub-environment terminates (“done”) when it either reaches the maximum episode length or the task is successfully completed. There are two common strategies for handling such terminations. One is the Fixed Episode Length mode, where environments are only reset after all sub-environments reach the maximum episode length. The other is the Partial Reset mode, where any sub-environment is reset immediately upon termination, without waiting for others. Figure 5(a) and Figure 5(b) illustrate the difference between these two modes. Within RLinf, to enable Fixed Episode Length, please set “ignore_terminations=True”, while to enable Partial Reset, please set “auto_reset=True” and “ignore_terminations=False”.
For PPO, users can flexibly choose between the two modes, which correspond to different optimization objectives. The Fixed Episode Length mode encourages the policy to maintain success until the end of an episode, aligning with the “success_at_end” objective defined in ManiSkill (Tao et al., 2025). In contrast, the Partial Reset mode emphasizes achieving success at least once during a rollout, i.e., the “success_once” objective. When the target is “success_once,” Partial Reset can substantially improve sample efficiency, as successful episodes are immediately reset and new trajectories are collected.
(a) Fixed Episode Length
(b) Partial Reset
(c) Valid Action Mask
Figure 5: Illustration of the three rollout modes. Section 4.2.1 and Section 4.2.2 provide detailed explanations of modes (b) and (c). In the RLinf-VLA framework, modes (a) and (b) are adopted for PPO, while modes (a) and (c) are adopted for GRPO.
Critic Design
To support PPO in VLA, the design of the critic is crucial. Using an entirely separate critic model would be computationally expensive and complicate GPU resource allocation, given the large size of VLA. To address this, we let the actor and critic share most of their parameters. Following the design in RL4VLA (Liu et al., 2025), we attach a lightweight value head (a three-layer MLP) on top of the language model component of VLA to estimate state values efficiently. Specifically, the value head takes the hidden representation $h^{0}$ from the final transformer block (Vaswani et al., 2023) at the position of the first action token and maps it to a scalar value prediction.
Value for Action Chunks
As discussed in Figure 4, two formulations of advantage estimation naturally lead to two corresponding formulations of value estimation for an action chunk $c_{t}=(a_{t,1},a_{t,2},\ldots,a_{t,C})$:
  1. 1.
    Chunk-level: The critic estimates the value of the entire chunk as a macro-action. Formally, $V:\mathcal{S}\to\mathbb{R}$.
  2. 2.
    Action-level: Each atomic action $a_{t,j}$ is assigned its own value estimate. Since only the initial observation $o_{t}$ is available at the beginning of the chunk, the critic outputs a $C$-dimensional vector, providing one value prediction per step within the chunk. Formally, $V:\mathcal{S}\to\mathbb{R}^{C}$.
Our framework supports both formulations, controlled by the parameter value_type. By default, the configuration sets value_type to match reward_type, ensuring consistency between reward aggregation and value estimation. We report results for both options using the OpenVLA-OFT model, and in general, the action-level formulation yields better performance.
4.2.2 GRPO
Group Design
The core of the GRPO algorithm lies in how groups are defined. In our framework, two trajectories are assigned to the same group if and only if they correspond to the same task and share the same initial state. This design mirrors the LLM setting, where multiple responses belong to the same group if and only if they are generated in response to the same prompt.
Valid Action Mask
During rollout, each environment runs for at most “max_episode_steps” steps. However, many tasks can be successfully completed earlier. This leads to two alternative strategies for policy optimization: (1) using the entire trajectory regardless of when success occurs, or (2) considering only the steps before task completion. Consistent with the Partial Reset mode discussed in Section 4.2.1, these correspond to two distinct optimization objectives, namely “success_at_end” and “success_once”. Our codebase supports both objectives. We refer to the latter as the Valid Action Mask setting, which can be enabled by setting “auto_reset=False” and “ignore_terminations=False”. Figure 5(c) visualizes this option. In experiments, we observed that under the GRPO setting, the valid action mask objective is generally more effective for improving policy performance. A detailed analysis is provided in the following section.
Loss Normalization by Trajectory Length
Unlike in LLM tasks, where longer responses are often more likely to be correct, successful trajectories in robotics are typically not longer than failed ones. To ensure that both successful and failed trajectories contribute equally during optimization, we normalize the policy loss by the trajectory length under the Valid Action Mask setting. Specifically, if a trajectory $\tau_{i}$ has $T_{i}^{\text{succ}}$ valid timesteps, the contribution of each timestep to the objective is scaled by $1/T_{i}^{\text{succ}}$. The GRPO objective under this setting is:
$J^{\text{GRPO}}(\theta)=\mathbb{E}_{o_{0}\sim D,\{\tau_{i}\}\sim\pi_{\theta_{\text{old}}}}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T_{i}^{\text{succ}}}\sum_{t=1}^{T_{i}^{\text{succ}}}\min\Big(\rho_{i,t}(\theta)\hat{A}_{i},\;\mathrm{clip}(\rho_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}\Big)\Bigg].$ (12)
This prevents longer trajectories from disproportionately dominating the gradient and ensures balanced learning across trajectories of varying lengths.
Success Rate Filter
Inspired by the dynamic sampling strategy in DAPO (Yu et al., 2025b), we introduce a success-rate filter for GRPO. This filter discards groups in which all trajectories either succeed or fail [^3] Following the practice in VeRL (Sheng et al., 2025), we set both lower and upper bounds on the average return of all trajectories within a group., since the advantage is non-zero only when both successful and failed trajectories coexist. Empirically, we find that this filtering mechanism accelerates convergence and consistently improves policy performance.
5 Experiment Results
In this section, we systematically investigate four key questions concerning the effectiveness of the proposed RLinf-VLA framework:
(1) Is RLinf-VLA high-performance? We evaluate RLinf-VLA on two representative testbeds, LIBERO and ManiSkill. The results demonstrate that RLinf-VLA achieves a success rate of 96.8% in large-scale mixed-task training involving up to 130 tasks, highlighting its strong capability to support large-scale multi-task learning.
(2) Is RLinf-VLA high-efficiency? We benchmark the framework across both GPU-parallelized and CPU-parallelized simulators, and observe that the optimal configuration differs across simulator types. This finding underscores the necessity of supporting diverse allocation modes. Notably, RLinf-VLA achieves up to $1.88\times$ speedup over existing frameworks.
(3) What are the best practices for applying PPO and GRPO to VLA training? Through extensive ablation studies, we identify the key factors that govern training performance. Moreover, we provide insights into the selection of common hyperparameters, offering practical guidelines for effectively deploying RL in VLA settings.
(4) Can models trained in simulation be deployed in the real world? To assess sim-to-real transfer, we conduct preliminary evaluations on six unseen object pick-and-place tasks. The results show that RL-trained models achieve higher zero-shot deployment success rates compared to SFT-trained models, demonstrating the promise of RL-based training for real-world applications.
Evaluation. Unless otherwise specified, all curves are smoothed using a Gaussian filter ($\sigma=1$), and success rate is computed under the “success_once” criterion, where an episode is considered successful if the success state is reached at least once.
5.1 High-performance
5.1.1 ManiSkill
Experiments Setup
  1. 1.
    Training tasks: “PutOnPlateInScene25Main-v3”. As previously mentioned, it is adopted in RL4VLA (Liu et al., 2025).
  2. 2.
    Evaluation tasks: Various out-of-distribution (OOD) test sets on vision, semantic, and position. Also adopted from RL4VLA. This part checks the capacity of VLA models on vision, language, and action, respectively. All models are evaluated under each sub-setting using 256 randomly episodes.
  3. 3.
    Hyperparameters: Please refer to YAML files in RLinf-VLA open-sourced code, including maniskill_ppo_openvla.yaml, maniskill_ppo_openvlaoft.yaml, maniskill_grpo_openvla.yaml, and maniskill_grpo_openvlaoft.yaml.
  4. 4.
    Base models: For the OpenVLA model, we adopt the pre-trained checkpoint available at HuggingFace [^4] https://huggingface.co/gen-robot/openvla-7b-rlvla-warmup. For the OpenVLA-OFT model, we perform our own LoRA fine-tuning using motion planning data collected from the “PutOnPlateInScene25Main-v3” task. The resulting LoRA model weights are also provided at HuggingFace [^5] https://huggingface.co/RLinf/RLinf-OpenVLAOFT-ManiSkill-Base-Lora.
Training Results
Figure 6 reports the training performance on ManiSkill across 25 tasks, using the OpenVLA and OpenVLA-OFT models trained with PPO and GRPO. Figure 6(a) shows the results for OpenVLA, while Figure 6(b) presents the OpenVLA-OFT results. Across all settings, reinforcement learning provides substantial performance gains, improving success rates by 45%–70% compared to the baseline. Notably, PPO consistently outperforms GRPO and exhibits greater stability for both OpenVLA and OpenVLA-OFT in the 25-task ManiSkill setting.
(a) OpenVLA
(b) OpenVLA-OFT
Figure 6: Training curves on ManiSkill “PutOnPlateInScene25Mani-v3” with OpenVLA and OpenVLA-OFT models, using PPO and GRPO algorithms. PPO consistently outperforms GRPO and exhibits greater stability.
Evaluation Results
Table 2 summarizes the evaluation results of RLinf-VLA models. For OpenVLA, we include four variants: the base model (“OpenVLA (Base)”), the RL4VLA baseline trained with PPO (“RL4VLA (PPO)”), and our RLinf-VLA models trained with GRPO (“OpenVLA (RLinf-GRPO)”) and PPO (“OpenVLA (RLinf-PPO)”). For OpenVLA-OFT, we report results on three variants: the LoRA fine-tuned base model (“OpenVLA-OFT (Base)”), along with the RLinf-VLA models trained using GRPO (“OpenVLA-OFT (RLinf-GRPO)”) and PPO (“OpenVLA-OFT (RLinf-PPO)”).
Although there is no fundamental algorithmic difference between the PPO-trained OpenVLA models in RLinf-VLA and in RL4VLA, our framework is carefully optimized to make better use of GPU resources. This enables training for more steps within a feasible time budget, leading to slightly better performance in RLinf-VLA compared to RL4VLA.
It is important to note that a direct comparison between OpenVLA and OpenVLA-OFT on OOD evaluation is not entirely fair, since their base models differ substantially in performance. Specifically, while OpenVLA improves from $39.10\%$ to $81.93\%$ after RL fine-tuning and OpenVLA-OFT improves from $18.29\%$ to $77.05\%$, the relative gain of OpenVLA-OFT is slightly higher despite its much weaker starting point. This indicates that the base model is very important and plays a crucial role in the final generalization.
Table 2: Evaluation results on ManiSkill. Values denote success rates.
In-Distribution Out-Of-Distribution
Vision Semantic Execution Avg.
OpenVLA (Base) 53.91% 38.75% 35.94% 42.11% 39.10%
RL4VLA (PPO) 93.75% 80.47% 75.00% 81.77% 79.15%
OpenVLA (RLinf-GRPO) 84.38% 74.69% 72.99% 77.86% 75.15%
OpenVLA (RLinf-PPO) 96.09% 82.03% 78.35% 85.42% 81.93%
OpenVLA-OFT (Base) 28.13% 27.73% 12.95% 11.72% 18.29%
OpenVLA-OFT (RLinf-GRPO) 94.14% 84.69% 45.54% 44.66% 60.64%
OpenVLA-OFT (RLinf-PPO) 97.66% 92.11% 64.84% 73.57% 77.05%
5.1.2 LIBERO
Experiments Setup
  1. 1.
    Tasks: We train a single unified model on all 130 tasks in LIBERO and evaluate its performance across the five LIBERO task suites: LIBERO-Object, LIBERO-Spatial, LIBERO-Goal, LIBERO-10, and LIBERO-90.
  2. 2.
    Hyperparameters: See the YAML configuration file in the RLinf-VLA open-source repository: libero_130_grpo_openvlaoft.yaml.
  3. 3.
    Base models: We fine-tune OpenVLA-OFT with LoRA ($r=32$) using demonstrations from all five LIBERO task suites: LIBERO-Object, LIBERO-Spatial, LIBERO-Goal, LIBERO-10, and LIBERO-90. The resulting LoRA model weights are also provided at HuggingFace  [^6] https://huggingface.co/RLinf/RLinf-OpenVLAOFT-LIBERO-130-Base-Lora.
Training Results
Figure 7 illustrates the training curve of OpenVLA-OFT with the GRPO algorithm across the five task groups. The x-axis represents the number of training epochs, while the y-axis indicates the corresponding success rate. The results show that the success rate improves substantially from approximately $73\%$ to $98\%$, yielding an overall performance gain of about $30\%$. These findings highlight the effectiveness of the GRPO design in RLinf-VLA for enhancing multi-task training.
Figure 7: Training curve of OpenVLA-OFT with GRPO on LIBERO-130.
Evaluation Results
Table 3 summarizes the evaluation results of the OpenVLA-OFT model trained on LIBERO-130 using RLinf-VLA across the five LIBERO task groups. For each task group, the success rate is computed as the average over 50 evaluation episodes per task. For instance, in the “LIBERO-90”, the reported success rate is averaged over a total of $90\times 50=4500$ episodes. The results show that the LIBERO-Spatial and LIBERO-Object group reaches over $99\%$, the LIBERO-Goal and LIBERO-90 task groups achieve up to $98\%$, and the LIBERO-10 task group exceeds $93\%$, yielding an overall average success rate of $98.11\%$. This corresponds to an average performance improvement of $32.68\%$ across task categories. Notably, our approach adopts a single model for all tasks, in contrast to the existing RL+VLA paradigm that trains and evaluates a separate model for each task group. This demonstrates RLinf-VLA’s capability to support large-scale multi-task RL.
Table 3: Evaluation results of the unified model on the five LIBERO task groups.
Spatial Object Goal 10 90 Avg.
OpenVLA-OFT (Base) 72.18% 71.48% 64.06% 48.44% 70.97% 65.43%
OpenVLA-OFT (RLinf-GRPO) 99.40% 99.80% 98.79% 93.95% 98.59% 98.11%
$\Delta$ Improvement +27.22{+27.22} +28.32{+28.32} +34.73{+34.73} +45.51{+45.51} +27.62{+27.62} +32.68{+32.68}
5.2 High-efficiency
To evaluate the efficiency of our framework, we conduct end-to-end experiments comparing throughput under different GPU allocation modes.
5.2.1 Experiment Setup
Evaluation Metric.
We use throughput as the evaluation metric for efficiency. Specifically, throughput is defined as the total number of rollout environment frames divided by the total wall-clock time of one training epoch, which approximately equals the sum of rollout time and training time. The evaluation is conducted on 8, 16, and 32 NVIDIA H100 (80GB) GPUs, distributed across 1, 2, and 4 nodes, respectively.
Tasks.
We evaluate our framework on both ManiSkill and LIBERO setups.
For ManiSkill, we select the “PutCarrotOnPlateInScene-v2” task for a quick test. It is adapted from “PutCarrotOnPlateInScene-v1” in ManiSkill (Tao et al., 2025), with modifications to the reset function. [^7] In the original reset implementation, additional simulation steps were applied to ensure that all objects remained static. However, this significantly increased the reset time compared to the step function and was incompatible with partial reset. In practice, we observed that these extra steps had a negligible impact on object states in the “PutCarrotOnPlateInScene” task, since the initial position of the carrot was already carefully calibrated. We therefore delete the additional simulation steps in the reset function. We use 256 parallel environments, each running for 80 steps.
For LIBERO, we adopt the LIBERO-10 task set for a quick test. The number of parallel environments is set to 64, 128, and 256 for 8-, 16-, and 32-GPU setups, respectively, with the corresponding number of environment steps set to 4096, 2048, and 1024. Since vectorized environments in LIBERO rely on multi-processing, the number of physical CPU cores per node becomes the upper bound for efficient rollout. Therefore, we scale the number of parallel environments in proportion to the number of nodes, and consequently to the total number of GPUs.
Baselines.
For ManiSkill, no existing framework supports multi-GPU rollout and training. Thus, we take the naive disaggregated allocation mode as the baseline, and compare it with our colocated mode and hybrid modes, where the hybrid mode includes both one-stage and two-stage fine-grained pipelining configurations.
For LIBERO, we compare against SimpleVLA-RL (Li et al., 2025), an open-source framework for VLA+RL training built on VeRL. Since SimpleVLA-RL only supports the colocated mode, we use it as the baseline and additionally evaluate our colocated mode and hybrid mode (with one-stage and two-stage fine-grained pipelining). We omit the disaggregated mode results for LIBERO, as during training the Training component never reuses the same GPUs as the rollout process. Consequently, even when accounting for the offload–onload overhead, the hybrid (one-stage) configuration provides superior performance to the disaggregated mode.
GPU Allocation Modes.
Tables 4 and 5 summarize the GPU allocation and offloading configurations used in our experiments.
Table 4: The setting of enable_offload among different GPU allocation modes.
Simulator Generation Training
Disaggregated False False False
Colocated True True True
Hybrid True True True
Table 5: Hyperparameters of GPU allocation modes.
# GPU Allocation Mode Simulator Generation Training
Disaggregated 0-1 2-3 4-7
8 GPUs Colocated 0-7 0-7 0-7
Hybrid 0-3 4-7 0-7
Disaggregated 0-3 4-7 8-15
16 GPUs Colocated 0-15 0-15 0-15
Hybrid 0-7 8-15 0-15
Disaggregated 0-7 8-15 16-31
32 GPUs Colocated 0-31 0-31 0-31
Hybrid 0-15 16-31 0-31
(a) GPU-parallelized Simulator ManiSkill with OpenVLA
(b) GPU-parallelized Simulator ManiSkill with OpenVLA-OFT
(c) CPU-parallelized Simulator LIBERO with OpenVLA-OFT
Figure 8: Throughput for different settings. “pipe” denotes the “pipeline_stage_num” for fine-grained pipelining.
5.2.2 Results
Figure 8 presents the end-to-end throughput of RLinf and baseline methods in VLA+RL training under different cluster sizes and placement strategies.
As shown in Figure 8(a), for the GPU-parallelized ManiSkill simulator with OpenVLA, RLinf-VLA achieves a $1.88\times$ speedup compared to the baseline disaggregated mode when using the hybrid configuration with pipe=2 on 8 GPUs. When scaling to 16 and 32 GPUs, additional overhead arises from model loading, offloading, and state switching. Consequently, the colocated mode generalizes less effectively than the disaggregated mode. However, the hybrid (pipe=2) configuration mitigates these costs through pipelined execution that overlaps computation and communication, still achieving a $1.61\times$$1.69\times$ speedup over the disaggregated mode. Furthermore, comparing hybrid configurations with pipe=1 and pipe=2, we observe that deeper pipelines yield better performance, suggesting that fine-grained pipelining effectively reduces GPU idle time (“bubbles”) during the rollout phase.
As shown in Figure 8(b), the results for OpenVLA-OFT exhibit substantial differences. In this case, the colocated mode and the hybrid (pipe=1) mode outperform the disaggregated mode and the hybrid (pipe=2) mode. This discrepancy stems from the use of action chunks. Since OpenVLA-OFT predicts an entire chunk of actions in each inference step while the simulator executes them sequentially, the simulator’s step execution time becomes much longer than the generation time (shifting from a 1:1 ratio to approximately 15:1). Consequently, the benefit of pipelining to hide generation latency diminishes. Moreover, since ManiSkill is GPU-parallelized and the rollout speed scales linearly with the number of parallel environments, allocating more GPU resources to the simulator makes the colocated and hybrid (pipe=1) modes more favorable.
Notably, when comparing the colocated mode and the hybrid (pipe=1) mode, we observe that the latter achieves progressively better throughput as the number of GPUs increases. This is because, in the colocated configuration, each GPU hosts both the simulator and the actor model simultaneously  [^8] Note that in the standard colocated setting, both the simulator and the actor model fully occupy all GPUs, and context switching is handled via onload and offload. However, since ManiSkill is GPU-parallelized, this approach incurs excessive overhead during multi-step interactions. Therefore, in our implementation, each GPU allocates half of its memory to each component.. In contrast, the hybrid configuration splits the cluster evenly, with each half of the GPUs dedicated to one component. As a result, the communication overhead in the colocated mode becomes increasingly non-negligible as the number of GPUs grows.
Figure 9: Latency breakdown for LIBERO with the OpenVLA-OFT setting.
For the CPU-parallelized LIBERO simulator with OpenVLA-OFT [^9] Previous results have shown that the disaggregated mode consistently underperforms the hybrid mode, as the training process in the disaggregated setup cannot utilize all GPUs. Therefore, the disaggregated mode is excluded from comparison here., the trend differs once again. As shown in Figure 8(c), the colocated mode achieves the best performance, yielding a $1.34\times$$2.27\times$ speedup over the SimpleVLA-RL implementation.
We further analyze the source of this speedup in Figure 9, where “Rollout” represents the total time for multi-step interactions between the generator and the simulator. Performance improvements are observed in both the rollout and training stages. Specifically, during rollout, RLinf-VLA employs a vectorized environment implementation that is more efficient than the hand-written multi-process environment workers used in SimpleVLA-RL. In addition, our implementation avoids redundant log-probability recomputation present in SimpleVLA-RL, further improving the rollout performance. During the training stage, RLinf adopts several system-level optimizations, including an adaptive communication mechanism, which leads to greater efficiency gains as the number of GPUs increases. Compared with the colocated mode, the hybrid mode incurs longer rollout times since LIBERO is a CPU-parallelized simulator. In this case, computation overlap provides little benefit, and allocating more compute resources to generation yields higher returns.
In summary, different types of simulators require distinct allocation modes, and varying interaction characteristics further necessitate adaptive configurations. RLinf-VLA supports three flexible allocation modes, enabling optimal system configurations under diverse scenarios.
5.3 Ablation Study
In this section, we present ablation studies under different setups and summarize key insights for PPO and GRPO training. In Section 5.3.1, we highlight PPO-specific strategies, such as step-level value estimation and partial reset. In Section 5.3.2, we discuss GRPO techniques, including trajectory length normalization, valid action masking, and success filtering. Finally, in Section 5.3.3, we analyze the effects of rollout data size and LoRA-based adaptations.
5.3.1 Tips for PPO
(a) Success rate in ManiSkill.
(b) Value loss in ManiSkill.
(c) Success rate in LIBERO-Goal
Figure 10: Ablation of value estimation type on PPO with action chunks.
(a) OpenVLA
(b) OpenVLA-OFT
Figure 11: Ablation study on partial reset with PPO in ManiSkill.
Action-level value estimation outperforms chunk-level estimation for PPO with action chunks.
Figure 10 shows the PPO training curves comparing action-level and chunk-level value estimation as described in Section 4.2.1. The x-axis denotes the number of training epochs, while the y-axis reports the success rate (Figure 10(a)) and value loss (Figure 10(b)) on the ManiSkill “PutOnPlateInScene25Main” task using the OpenVLA-OFT model. Step-level estimation consistently yields higher success rates and lower value loss, demonstrating more effective learning and faster policy improvement. This advantage is not task-specific: similar results are observed in the LIBERO-Goal benchmark (Figure 10(c)).
Partial reset substantially improves sample efficiency.
Figure 11 presents the PPO training curves for the Partial Reset and Fixed Episode Length rollout modes discussed in Section 4.2.1. The x-axis indicates the number of training epochs, and the y-axis shows the success rate. Since the optimization objective is “success_once,” Partial Reset leads to a significantly higher success rate. For a given number of training epochs, the success rate under Partial Reset consistently exceeds that of the Fixed Episode Length mode. This trend is observed regardless of the model type (OpenVLA or OpenVLA-OFT).
5.3.2 Tips for GRPO
Trajectory length normalization in GRPO.
Figure 12: Ablation study on trajectory length normalization.
As discussed in Section 4.2.2, normalizing the loss by trajectory length aims to reduce bias when episodes vary in length. Figure 12 reports results on LIBERO-Goal, where the x-axis denotes the number of training epochs and the y-axis the success rate. The results show that incorporating trajectory length normalization (“w/ Norm”) can lead to substantially higher performance compared with the unnormalized setting (“w/o Norm”).
Valid action mask in GRPO.
We also study the effect of valid action masking, introduced in Section 4.2.2. Since the optimization objective is defined with respect to “success_once”, excluding actions after reaching the success state improves sample efficiency and avoids redundant updates. Moreover, applying the valid action mask naturally results in shorter trajectories, which further benefits trajectory length normalization. Figure 13(a) shows results on LIBERO-Goal. Comparing the “w/o Mask, w/o Norm” and “w/ Mask, w/o Norm” curves, the setting with a valid action mask achieves consistently better performance. The “w/ Mask, w/ Norm” curve demonstrates that combining the mask with trajectory length normalization provides an additional improvement.
(a) LIBERO-Goal
(b) ManiSkill
Figure 13: Ablation studies on valid action mask in GRPO.
However, the effect of these techniques is task-dependent. In the ManiSkill setting, we do not observe clear benefits from either valid action masking or trajectory length normalization. As shown in Figure 13(b), the performance of the “w/ Mask, w/ Norm” configuration is comparable to that of the “w/o Mask, w/o Norm” baseline, suggesting that the effectiveness of these strategies may vary across environments.
(a) OpenVLA, ManiSkill
(b) OpenVLA-OFT, ManiSkill
(c) OpenVLA-OFT, LIBERO-Goal
Figure 14: Ablation studies on success rate filtering under different settings.
Success rate filtering can improve training stability in some settings of GRPO.
As discussed in Section 4.2.2, we implement a success rate filter for GRPO that discards groups in which all trajectories have identical cumulative rewards. This mechanism can improve the stability of GRPO training. For instance, in the OpenVLA ManiSkill setting, Figure 14(a) shows that training without the filter (“w/o Filter”) exhibits a clear collapse around step 400, whereas enabling the filter (“w/ Filter”) largely alleviates this issue. However, the benefit of the filter is not universal: in the OpenVLA-OFT ManiSkill setting (Figure 14(b)) and the OpenVLA-OFT LIBERO-Goal setting (Figure 14(c)), its effectiveness is much less pronounced.
5.3.3 Hyperparameters Choice
Effect of rollout data size.
The rollout batch size has a notable influence on RL performance, especially for on-policy algorithms such as PPO and GRPO. In general, larger rollouts per epoch enable more substantial policy improvement within each training iteration. As illustrated in Figure 15, larger rollouts consistently achieve higher success rates when evaluated by training epochs.
(a) PPO, LIBERO-10
(b) GRPO, LIBERO-130
Figure 15: Ablation study on rollout data size. Darker colors correspond to larger rollout datasets. “x Trajs” denotes the number of trajectories included in the rollout data per iteration.
Using LoRA may not directly affect performance but often requires different hyperparameters.
Figure 16 shows the training curves with and without LoRA. The x-axis denotes the number of training epochs, while the y-axis shows the success rate in ManiSkill. The curves are overall similar, suggesting that LoRA itself does not substantially change performance. However, the choice of whether to use LoRA can influence the optimal hyperparameters. For example, in GRPO experiments (Figure 16(b)), using the same learning rate of $1\times 10^{-4}$ leads to normal improvement with LoRA, but the success rate without LoRA collapses to zero. In contrast, when the learning rate is reduced to $1\times 10^{-5}$, the non-LoRA setting also achieves stable improvement. These results indicate that different LoRA configurations may require separate hyperparameter tuning.
(a) PPO
(b) GRPO
Figure 16: Ablation of LoRA in ManiSKill with OpenVLA-OFT model.
5.4 Real-World Deployment
We further evaluate the transferability of simulation-trained models by deploying them to the real world.
Experimental Setup
The hardware setup consists of a Franka Panda robotic arm and an Intel RealSense D435 camera. The relative pose between the camera and the robot is calibrated to ensure consistency with the simulation environment. For real-world evaluation, we compare the “OpenVLA (RLinf-PPO)” model with “OpenVLA (SFT)”. For “OpenVLA (SFT)”, we train OpenVLA to converge on datasets ranging from a few hundred to 64k expert trajectories (1.26M transitions) in simulation. We observe that performance plateaus at roughly 16k trajectories, so we adopt the 16k-trajectory SFT checkpoint as the baseline for comparing RL fine-tuning methods. The evaluation involves six previously unseen objects (pepper, cucumber, banana, kiwi fruit, mangosteen, and sponge), each tasked with being placed into a bowl. For every object, five trials are conducted under different initial configurations of both the object and the bowl.
Table 6: Real-world deployment results. “Pick” indicates that the policy successfully grasped the object, while “Success” indicates that the object was both picked and placed into the bowl.
Object Type OpenVLA (SFT) OpenVLA (RLinf-PPO)
Pick Success Pick Success
Pepper 1/5 0/5 2/5 1/5
Cucumber 1/5 0/5 1/5 1/5
Banana 0/5 0/5 3/5 3/5
Kiwi 0/5 0/5 2/5 1/5
Mangosteen 1/5 0/5 1/5 1/5
Sponge 0/5 0/5 4/5 1/5
Total 3/30 0/30 13/30 8/30
Figure 17: Example of real-world deployment with an RL-trained model performing the task of placing a banana into a bowl.
Results
Table 6 reports the average performance in real-world trials. OpenVLA (RLinf-PPO) substantially outperforms OpenVLA (SFT). Figure 17 illustrates a representative example, where the model successfully places a banana into a bowl. Notably, beyond basic calibration between simulation and the real world, no additional sim-to-real adaptation techniques were employed. Despite the preliminary nature of the experiment, the RL-trained policy demonstrates zero-shot transfer to real-world tasks, revealing the strong generalization capability enabled by simulation-based reinforcement learning. We observe that the SFT policy tends to exhibit excessive motion and overshooting, often resulting in inaccurate grasp poses and task failures. In contrast, the RL-trained policy, though showing minor oscillations, iteratively refines the end-effector pose and achieves a higher grasp success rate.
6 Conclusion
In this work, we introduced RLinf-VLA, a unified and efficient framework for reinforcement learning-based training of Vision-Language-Action models. RLinf-VLA integrates multiple simulators, algorithms, and VLA architectures, while providing flexible execution modes and system-level optimizations that significantly improve training efficiency. Extensive experiments demonstrate that a single unified model achieves state-of-the-art success across a wide range of simulated tasks, and RL-trained policies exhibit stronger zero-shot generalization on a real-world Franka robot compared to supervised fine-tuning. Moreover, our study distills actionable best practices for both PPO and GRPO, guiding future research in RL-based VLA training. By open-sourcing RLinf-VLA with ongoing maintenance, we provide the community with a foundation to accelerate, standardize, and scale research in embodied intelligence.
7 Future Work
Our long-term goal is to advance the capabilities of VLA models through reinforcement learning. Achieving this requires a reliable and extensible infrastructure, which we will continue to actively maintain and expand. Below, we outline several directions for future development.
  • We plan to broaden simulator support. Integration with RoboTwin (Mu et al., 2025) and IsaacLab (Developers, ) is currently in progress and will be released soon, enabling researchers to explore a wider range of embodied environments.
  • Second, we will extend the set of supported models. In particular, integrations with $\pi_{0}$ (Black et al., 2024) and $\pi_{0.5}$ (Intelligence et al., 2025) will be available in the near future, along with algorithmic tuning to ensure strong performance out of the box.
  • Third, we aim to enrich the algorithmic coverage of the framework. Beyond the current on-policy methods, we are working to incorporate classical off-policy algorithms such as Soft Actor-Critic (SAC) (Haarnoja et al., 2018), which are known for their sample efficiency and stability.
  • Finally, while our current focus has been on simulation-based RL, we will also integrate real-world reinforcement learning into the codebase. This step is critical for bridging the gap between simulation and real-world deployment, and will further highlight the practicality of our framework.
In the longer term, we envision this framework as a foundation for establishing standardized benchmarks and fostering community collaboration, thereby accelerating progress in the development and deployment of VLA models.